Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
vardaro, zendri - 18:05 Monday 19 January 2026 (68521) Print this report
Optimization of alignment with EQB1 delay line

The goal of the shift was to measure the amount of losses and phase noise during a CC phase scan with the squeeed beam folded into the EQB1 bench (delay line configuration). 
We worked in stand alone configuration, i.e. with only the CC PLL locked.

First part of the shift we struggled because the magnitude of the 4MHz beat note on the HD detector is lower of what we had in the past (5.3mV vs 7.5 mV of last Thursday) but the amount of squeezing generated by the OPA is only sligthly decreased from 8.8dB of last week to 8.6 dB of today.

We checked also with the retroreflector in order to see if was a clipping inside the delay line but no improvement on that side. We decided to go ahead in any case and we started by optimizing the alignment.

1) checking the level of squeezing we moved the HD PDs in X an Y trying to optimize it (5steps each movement)

2) Checking the resonance of mirror mounts in the  HD DIFF AUDIO spectrum we optimized the EQB1 FI HWP and  we managed to kill al of them

3) Adding scan on EQB1 HD M6 mirror X and Y DOF with ramps we managed to optimize the Delay alignment using DL M6 and M5 mirrors (20-30 steps per time)
4) we checked the level of sqz and asqz in this configuration and we managed to obtain about 8.1dB of ASQZ and 6.2 dB of SQZ

we run two phase scan of the CC beam and we are analyzing the data.

Te 4 MHz magnitud is less stable than usual. Tomorrow we will start from this in checking the system

Images attached to this report
Comments to this report:
vardaro - 13:50 Monday 26 January 2026 (68568) Print this report

We performed further analysis on this topic.

- We noticed that despite the (anti)squeezing level was not changing, the mag passed from 5.5mV to 8mV and viceversa. See fig 1 and 5. We noticed that this is correlated to the PumpPhase PD DC but not with the power of the SQZ Main Laser

- We managed to do a CC phase scan with HD detector in both the states (Fig3) and we managed to fit both of them Fig 3 and 4

With the fit we found 10.3% of losses with 40mrad o phase noise. To be compared with the 11% of losses and 40mrad of phase noise found in september 2022. Note that we used a completely different investigation method that we are still checking. So the increase phase noise can be also due to the different measurement method.
Next steps is to validate these measurement, extend the measurements up to the retroreflector and to unnderstand the cause in the jumps of the 4 MHz magnitude 

Images attached to this comment
vardaro, zendri, zambon, kranzhoff - 17:52 Wednesday 04 February 2026 (68620) Print this report

02/02/2025 Improvement of delay line alignment

We started the work with the laser in the good state, High Magnitude ~8mV so we decided to profit witht he sistuation and work with the delay line alignment.

We acted mostly on DL_M4 in both the degree of freedom and we managed to improve the magnitude level and the antisqueezing level. At the end of the operation we also checked how much the squeezing was improved and we saw an improvement of ~0.3dB

We performed a Coherent control loop phase scan and we managed to measure ~8dB of anti squeezing and 6.3 dB of squeezing with 8.3 dB of generated squeezing. (Fig1). Phase noise 63mrad losses 8%

03/02/2025 CC Laser mode hopping

We started the activity in the bad state of the CC laser. So we decided to investigate about the cause of this issue. For us it is impossible to improve the alignment if the amplitude of the 4 MHz magnitude on the HD detector is fluctuating.

The first test was to change the gain of the CC PLL to see if it was a servo bump of the loop causing some issue. But we did not see any effect in the magnitude

As second step we tried to keep again the MAIN PLL unlocked and we chenged the temperature of the main Laser with the CC PLL locked. We increased the absolute number of the Temperature set. We started from 31568 as starting value for the slow output of the CC PLL and the CC PLL started to unlock fbetween 31720 and 31740, i.e. when the CC Laser temperature was changed by 40mK, i.e. its frequency by about 120 MHz. In this situation we did not manage to lock anymore the CC PLL. During this change of temperature we noticed that the HD 4MHz magnitude decreased every step. Possible explanation: we started with the laser in a multimode configuration and after 120MHz the energy was more in the other peak and the CC PLL was not able to lock anymore.

We wanted to explore the other direction to understand if the laser was more stable and to understan how much we need to move the CC PLL frequency  and the MAIN Laser frequency. At that point Fiodor pointed out this entry https://logbook.virgo-gw.eu/virgo/?r=51719 in wich they experienced the same situation.

The idea is to add EQB1_HD_MIR i.e a mirror in front of the HD detector to redirect the beam going to PD2 toward a photodiode and two cameras out of the HD box. The name of the PD channel is EQB1_IR_PD_MONI. HD_MIR pos = 0 means that the Mirror is out from the way, and HD_MIR pos = +9000 steps means that the mirror is in the beam and the PD receive light. After that we remembered that the CC power is very low and it is impossible to see on the photodiode without removing the filter in front of it. So we removed the filter in front the external PD that is called "IR PD Filter" in the process EQB1_ThorAct.

We then scanned the OPA cavity, after OPA cavity in the direction of IR_PD_MONI, the Green beam is filtered and only the LO and CC beam are sent toward this PD. We thus shuttered LO and on the PD we saw only the CC. Looking the cavity scan (we saw multiple peaks) and moving the CC PLL by -1400 steps only one peak was visible (Fig2).

1400steps means 0.35V i.e. 0.35C in temperature of the laser. 

We wanted to test the system in this point so we moved the temperature of all the laser. So the new working point is: PLL_MAIN = 30120 CC_PLL=30038 SC_PLL=37966.  The starting point was: MAIN_PLL= 31544 CC_PLL=31423 SC_PLL = 38040.

In this point we bot checked the the stability with the system in anti-squeezing and we also di another phase scan of the CC. See fig 3 and 4. Phase noise 40mrad losses 10%

 

In this configutation the system is more stable but now squeezing can not be injected into the interferometer because the MAIN LASER of the squeezer is about 1.2GHz faar from Virgo Laser.

This problem started after this work in last july https://logbook.virgo-gw.eu/virgo/?r=67348 in wich the temperature of the ITF laser was moved by 1.4degree from 22.7 to 24.1C. In order to lock the MAIN PLL in this new working point the solution should be move the PSL Laser by ~0.35 degree i.e around 23.75 C.

  

03/02/2025 Check During night

We checked the behaviour during the night and we noticed that at the beginning was not aligned. After that was stable for a while and that some multimode behaviout occurred again. So probably 0.35 C is not enough and it is better to  change by 0.45 i.e to 23.65 C if it is possible from INJ system 

Images attached to this comment
vardaro - 18:05 Wednesday 04 February 2026 (68631) Print this report

We noticed that in all the scans the shot noise level was not well normalized and was aeoind -0.1dB. So we changed the normalization factor from 276548 to 278820

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×